# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ Appendix F Oracle Inequality
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widetilde{\pi}_{T},{\widehat{p}}^{*}_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+n^{\frac{3d^{*}+13}{12(d^{*}+3)}}m^{-\frac{1}{4}}\right).$
Moreover, if $m>n^{\frac{d^{*}+5}{d^{*}+3}}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widetilde{\pi}_{T},{\widehat{p}}^{*}_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}\right).$
Appendix G Main Result
Proof of Theorem 4.12.
In our framework, we have the following total error decomposition:
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T},p_{data})]$
$\leq\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T},(\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})_{\#}p_{data})]+\mathbb{E}_{\mathcal{Y}}[W_{2}((\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})_{\#}p_{data},(\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})_{\#}\widetilde{p}_{data})]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}_{\mathcal{Y}}[W_{2}((\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})_{\#}\widetilde{p}_{data},\widetilde{p}_{data})]+W_{2}(\widetilde{p}_{data},{p_{data}})$
$\leq\gamma_{\boldsymbol{D}}\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widetilde{\pi}_{T},\widehat{p}^{*}_{data})]+(\gamma_{\boldsymbol{D}}\gamma_{\boldsymbol{E}}+1)W_{2}(\widetilde{p}_{data},p_{data})+\mathbb{E}_{\mathcal{Y}}[\mathcal{H}(\widehat{\boldsymbol{\boldsymbol{E}}},\widehat{\boldsymbol{D}})]^{1/2},$
where the second inequality follows from
$\mathbb{E}_{\mathcal{Y}}[W_{2}((\widehat{\boldsymbol{D}}\circ\widehat{\boldsymbol{E}})_{\#}\widetilde{p}_{data},\widetilde{p}_{data})]\leq\mathbb{E}_{\mathcal{Y}}[\mathcal{H}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})^{1/2}]\leq\mathbb{E}_{\mathcal{Y}}[\mathcal{H}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})]^{1/2}.$
Therefore, under Assumptions 4.3-4.5, we obtain
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T},p_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+n^{\frac{3d^{*}+13}{12(d^{*}+3)}}m^{-\frac{1}{4}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\mathcal{M}^{-\frac{1}{2(d+2)}}+\sqrt{\delta_{0}}\right).$
Moreover, if $m>n^{\frac{d^{*}+5}{d^{*}+3}}$ and $\mathcal{M}>n^{\frac{d+2}{3(d^{*}+3)}}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T},p_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\sqrt{\delta_{0}}\right).$

### v2

@@ D.1 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})]$
$\left|\mathrm{(I)}-\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\int_{0}^{T}\mathbb{E}_{\mathbb{P}}\|\widehat{\mathbf{s}}(1-t,\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right]\right|\leq\epsilon.$
Taking $\epsilon=n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n$, then there exists $\Delta_{n}>0$ such that if $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})\leq\Delta_{n}$, then
$(\mathrm{I})\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n.$
Next, we bound the second term $\mathrm{(II)}$. For any $\epsilon>0$, there exists a constant $C>0$ such that
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}$
$\lesssim\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t_{i}}\|_{\infty}>1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}}\}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}>1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t_{i}}\|_{\infty}\leq 1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}},\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\lesssim\frac{\log n}{1-t_{i}}\mathbb{P}\left(\|\mathbf{x}_{t_{i}}\|_{\infty}>1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}}\right)+\frac{\log n}{1-t_{i}}\mathbb{P}\left(\|\mathbf{x}_{t}\|_{\infty}>1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t_{i}}\|_{\infty}\leq 1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}},\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}.$
By Lemma F.4, we have
$\mathbb{P}\left(\|\mathbf{x}_{t_{i}}\|_{\infty}>1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}}\right)\lesssim\sigma_{1-t_{i}}\epsilon,\leavevmode\nobreak\ \mathbb{P}\left(\|\mathbf{x}_{t}\|_{\infty}>1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\right)\lesssim\sigma_{1-t}\epsilon.$
According to Lagrange’s Mean Value Theorem and Lemma F.3, there exists $\mathbf{x}^{\prime}$ between $\mathbf{x}_{t_{i}}$ and $\mathbf{x}_{t}$ such that
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|\mathbf{I}_{\{\|\mathbf{x}_{t_{i}}\|_{\infty}\leq 1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}},\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$=\|\partial_{\mathbf{x}}\nabla\log q_{1-t_{i}}(\mathbf{x}^{\prime})\|\|\mathbf{x}_{t_{i}}-\mathbf{x}_{t}\|\mathbf{I}_{\{\|\mathbf{x}_{t_{i}}\|_{\infty}\leq 1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}},\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\lesssim\frac{\log\epsilon^{-1}}{\sigma^{2}_{1-t_{i}}}\cdot\|\mathbf{x}_{t_{i}}-\mathbf{x}_{t}\|.$
Then, we have
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t_{i}}\|_{\infty}\leq 1+C\sigma_{1-t_{i}}\sqrt{\log\epsilon^{-1}},\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\lesssim\frac{\log^{2}\epsilon^{-1}}{\sigma^{4}_{1-t_{i}}}\mathbb{E}_{\mathbb{P}}\|\mathbf{x}_{t_{i}}-\mathbf{x}_{t}\|^{2}\lesssim\frac{\log^{2}\epsilon^{-1}}{(1-t_{i})^{2}}\cdot(t-t_{i}).$
Therefore, we obtain
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})\|^{2}\mathrm{d}t$
$\lesssim\sum_{i=0}^{K-1}\left[\frac{(t_{i+1}-t_{i})\epsilon\log n}{\sqrt{1-t_{i}}}+\frac{(t_{i+1}-t_{i})\epsilon\log n\sqrt{1-t_{i}}}{1-t_{i}}+\frac{\log^{2}\epsilon^{-1}}{(1-t_{i})^{2}}\cdot(t_{i+1}-t_{i})^{2}\right]$
$\lesssim\frac{T\epsilon\log n}{\sqrt{1-T}}+\frac{T\log^{2}\epsilon^{-1}}{(1-T)^{2}}\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})$
$\lesssim\epsilon n^{\frac{\beta}{d^{*}+2\beta}}\log n+n^{\frac{4\beta}{d^{*}+2\beta}}\log^{2}\epsilon^{-1}\max_{0\leq i\leq K-1}(t_{i+1}-t_{i}).$
Taking $\epsilon=n^{-\frac{3\beta}{d^{*}+2\beta}}$ and $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})=\mathcal{O}(n^{-\frac{6\beta}{d^{*}+2\beta}})$, we obtain
$\mathrm{(II)}\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{2}n.$
Finally, we bound the last term $\mathrm{(III)}$. By Lemma F.4, for any $\epsilon>0$, there exists a constant $C>0$ such that
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}$
$\lesssim\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}>1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\lesssim\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{\log n}{1-t_{i}}+\frac{\log n}{1-t}\right)\mathbb{P}\left(\|\mathbf{x}_{t}\|_{\infty}>1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\right)$
$\lesssim\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{\log n}{1-t_{i}}+\frac{\log n}{1-t}\right)\sqrt{1-t}\cdot\epsilon.$
According to Lagrange’s Mean Value Theorem and Lemma F.3, there exists $t^{\prime}$ between $t_{i}$ and $t$ such that
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}$
$=\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t^{\prime}}(\mathbf{x}_{t})\mathbf{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq 1+C\sigma_{1-t}\sqrt{\log\epsilon^{-1}}\}}(\mathbf{x}_{t})\|\cdot(t-t_{i})^{2}$
$\lesssim\frac{1}{(1-t^{\prime})^{\frac{3}{2}}}\cdot\log^{\frac{3}{2}}\epsilon^{-1}\cdot(t-t_{i})^{2}$
$\lesssim\frac{1}{(1-T)^{\frac{3}{2}}}\cdot\log^{\frac{3}{2}}\epsilon^{-1}\cdot(t-t_{i})^{2}.$
Therefore, we have
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbb{P}}\|\nabla\log q_{1-t_{i}}(\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t$